OpenStudy (anonymous):

can you factor this? 32x^3-1372x^3y^3

OpenStudy (anonymous):

The Wolfram factorization is an answer. If you had to do this manually, you might be able to discover y = 2/7 but not easily, so I would just stop at 4x^3(8-343y^2) in that case.

OpenStudy (anonymous):

\[=4x^3(8-343y^3)\] \[=4x^3(2^3-(7y)^3)\] then \(2^3-(7y)^3 can be factorised identity(a^3-b^3)=(a-b) (a^2+a b+b^2)